Compton Meeting, January 11, 2012 (F226)

From PolWiki
Jump to: navigation, search



  1. Laser repair update (Gaskell)
  2. Entrance function status and laser polarization (Dalton)

Photon Detector

  1. LED linearity system (Cornejo)

Electron Detector

  1. Detector commissioning studies: what remains? (Narayan, Dutta)
  2. Parallel root analysis (Narayan)
  3. Special studies: beam position scan from Dec., rate dependence

Talks for Hall C Meeting

  1. Qweak polarimetry (Tvaskis) [1]
  2. Hall C polarimetry at 12 GeV (Gaskell) Draft


  1. Switching to the split backplane ELMA Crate ?


Attending:D. Gaskell, M. Dalton, D. Jones, V. Tvaskis, A. Asaturyan, H. Mkrtchyan

By phone: Jeff Martin


  1. Laser in hall repaired - some noise issues initially that mysteriously went away Sunday. 2nd laser needs new diode - waiting for quote.
  2. Entrance/transfer function: Mark is working on fits to entrance function. Also, taking data in order to apply similar technique to exit transfer function.

Photon Detector:

  1. Reviewed asymmetries from Don's log entry ( since Hamlet had not yet seen them.
  2. Discussed LED system status - Don indicated that in conversation with Juan Carlos, he heard there might be some cross-talk between the LEDs.
  3. May schedule a dedicated photon detector meeting to see where things stand, and make a plan going forward.

Electron detector:

  1. Reviewed Dipangkar's latest log entry ( Edet rates vary in a manner similar to photon detector. However, 3/4 rates show less time dependence. Indication of rate-dependent inefficiency?
  2. Dave thinks it's important to make a plan to study the dead time.
  3. Vladas notes some concern about charge normalization - Dave will look at BCM calibration data from last fall.
  4. Dave asked Don to make a run list for the position scan studies from December - then we'll get "best" edet polarization values from Vladas to look at correlation.


  1. Swapping in ELMA crate? Defer until more discussion with Amrendra.